OpenStudy (anonymous):

plse solve it.. its tan(x-pie/2)*cos(3pi/2+x)-sin^3(7pi/2-x)/cos(x-pi/2)*tan(3pi/2+x) simplifies to a.(1+cos^2x) b.sin^2x c.-(1+cos^2x) d.cos^2x

OpenStudy (anonymous):

plseee helppp

OpenStudy (adunb8):

try wolfram alpha xp

OpenStudy (anonymous):

@adunb8 who is wolfram alpha xp

OpenStudy (anonymous):

\[\tan(x-\pi/2)*\cos(3\pi/2+x)-\sin^3(7\pi/2-x)/\cos(x-\pi/2)*\tan(3\pi/2=x)\]

OpenStudy (anonymous):

Have they given you the formula to find sin^3(7pi/2-x)???

OpenStudy (amoodarya):

sin(7pi/2-x) = sin(7pi/2) cos(x) - cos(7pi/2) sin(x)